In 30 years, you plan to set up a fellowship fund for your university that pays out $100,000/year in perpetuity with an annually compounded discount rate of 5%. In order to set up the fund in 30 years, how much do you need to save each year (starting this year) assuming you can get a semi-annually compounded return of 10% on your savings for the next 30 years?
1
Expert's answer
2019-10-31T10:41:18-0400

The formula for perpetuity is: PV = CF/r = 100,000/0.05 = 2,000,000.

To get this sum in 30 years we should use the formula for future value of annuity:

"P =FV\u00d7r\/((1 + r)^n - 1)"

P = 2,000,000×0.1/(1.1^30 - 1) = 12,158.5 is the annual payment.